Papers by Andrew C. Haas

1. “Search for Randall-Sundrum gravitons with 1 fb-1 of data from ppbar collisions at sqrt(s)=1.96 TeV” V. M. Abazov et al. [D0 Collaboration] arXiv:0710.3338 [hep-ex] FERMILAB-PUB-07-553-E(2007)

2. “Measurement of the pp̄ → W Z + X cross-section at \sqrt{s} = 1.96-TeV and limits on WWZ trilinear gauge couplings” V. M. Abazov et al. [D0 Collaboration] arXiv:0709.2917 [hep-ex] FERMILAB-PUB-07-488-E(2007) (Submitted to Phys.Rev.D)

3. “Measurement of the muon charge asymmetry from W boson decays”

V. M. Abazov et al. [D0 Collaboration] arXiv:0709.4254 [hep-ex] FERMILAB-PUB-07-493-E(2007) (Submitted to Phys.Rev.D)

4. “Search for flavor-changing-neutral-current D meson decays”

V. M. Abazov et al. [D0 Collaboration] arXiv:0708.2094 [hep-ex] FERMILAB-PUB-07-423-E(2007) (Submitted to Phys.Rev.Lett.)

5. “Search for the lightest scalar top quark in events with two leptons in pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] arXiv:0707.2864 [hep-ex] FERMILAB-PUB-07-352-E(2007) (Submitted to Phys.Lett.B)

6. “Search for B_{s} → μ^{+}μ^{−} at D0”

V. M. Abazov et al. [D0 Collaboration] arXiv:0707.3997 [hep-ex] FERMILAB-PUB-07-395-E(2007)

7. “Measurement of the tt̄ production cross-section in pp̄ collisions using dilepton events” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 76, 052006 (2007) [arXiv:0706.0458 [hep-ex]]

8. “Direct observation of the strange b baryon Xi_{b}^{−}”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 99, 052001 (2007) [arXiv:0706.1690 [hep-ex]]

9. “Measurement of the Lambda_{b}^{0} lifetime using semileptonic decays”

V. M. Abazov et al. [D0 Collaboration] arXiv:0706.2358 [hep-ex] FERMILAB-PUB-07-205-E(2007) (Submitted to Phys.Rev.Lett.)

10. “Search for stopped gluinos from pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 99, 131801 (2007) [arXiv:0705.0306 [hep-ex]] 11. “Zγ production and limits on anomalous ZZγ and Zγγ couplings in pp̄ collisions at \sqrt{s} =

1.96- TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 653, 378 (2007) [arXiv:0705.1550 [hep-ex]]

12. “Search for a Higgs boson produced in association with a Z boson in pp̄ collisions” V. M. Abazov et al. [D0 Collaboration] arXiv:0704.2000 [hep-ex] FERMILAB-PUB-07-076-E(2007) (Submitted to Phys.Lett.B.)

13. “Measurement of the Λ_{b} lifetime in the exclusive decay Λ_{b} → J/ψΛ”

V. M. Abazov et al. [D0 Collaboration] arXiv:0704.3909 [hep-ex] FERMILAB-PUB-07-094-E(2007) (Submitted to Phys.Rev.Lett.)

14. “Search for production of single top quarks via flavor-changing neutral currents at the Tevatron” V. M. Abazov et al. [D0 Collaboration] arXiv:hep-ex/0702005 FERMILAB-PUB-07-031-E(2007) (Submitted to Phys.Rev.Lett.)

15. “Measurement of the top quark mass in the lepton + jets channel using the ideogram method” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 75, 092001 (2007) [arXiv:hep-ex/0702018]

16. “Measurement of the shape of the boson rapidity distribution for pp̄ → Z/gamma^{∗} → e^{+}e^{−} + X events produced at \sqrt{s} of 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 76, 012003 (2007) [arXiv:hep-ex/0702025]

17. “Combined D0 measurements constraining the CP-violating phase and width difference in the B/s0 system” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 76, 057101 (2007) [arXiv:hep-ex/0702030]

18. “Measurement of the branching fraction Br(B^{0}(s) → D_{s}^{(∗)}D_{s}^{(∗)})”

V. M. Abazov et al. [D0 Collaboration] arXiv:hep-ex/0702049 FERMILAB-PUB-07-047-E(2007) (Submitted to Phys.Rev.Lett.)

19. “Measurement of the charge asymmetry in semileptonic B_{s} decays”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 98, 151801 (2007) [arXiv:hep-ex/0701007]

20. “Lifetime difference and CP-violating phase in the B_{s}^{0} system”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 98, 121801 (2007) [arXiv:hep-ex/0701012]

21. “Search for single production of scalar leptoquarks in p anti-p collisions decaying into muons and quarks with the D0 detector” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 647, 74 (2007) [arXiv:hep-ex/0612012]

22. “Search for techniparticles in e + jets events at D0”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 98, 221801 (2007) [arXiv:hep-ex/0612013]

23. “Measurement of the pp̄ → tt̄ production cross section at \sqrt{s} = 1.96-TeV in the fully hadronic decay channel” V. M. Abazov et al. [D0 Collaboration] arXiv:hep-ex/0612040 FERMILAB-PUB-06-426-E(2006) (Submitted to Phys.Rev.D)

24. “Evidence for production of single top quarks and first direct measurement of —V(tb)—” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 98, 181802 (2007) [arXiv:hep-ex/0612052]

25. “Measurement of the t anti-t production cross section in p anti-p collisions at s**(1/2) = 1.96-TeV using secondary vertex b tagging” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 112004 (2006) [arXiv:hep-ex/0611002]

26. “Search for the pair production of scalar top quarks in the acoplanar charm jet topology in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 645, 119 (2007) [arXiv:hep-ex/0611003]

27. “Measurement of the CP-violation parameter of B0 mixing and decay with p anti-p –¿ mu mu X data” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 092001 (2006) [arXiv:hep-ex/0609014]

28. “Measurement of B_{d} mixing using opposite-side flavor tagging”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 112002 (2006) [arXiv:hep-ex/0609034]

29. “Measurement of the W boson helicity in top quark decay at D0”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 75, 031102 (2007) [arXiv:hep-ex/0609045]

30. “Measurement of the top quark mass in the lepton + jets final state with the matrix element method” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 092005 (2006) [arXiv:hep-ex/0609053]

31. “Measurement of the top quark mass in the dilepton channel”

V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 655, 7 (2007) [arXiv:hep-ex/0609056]

32. “Limits on anomalous trilinear gauge couplings from W W –¿ e+ e-, W W –¿ e+- mu-+, and W W –¿ mu+ mu- events from p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 057101 (2006) [Erratum-ibid. D 74, 059904 (2006)] [arXiv:hep-ex/0608011]

33. “Search for pair production of scalar bottom quarks in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 171806 (2006) [arXiv:hep-ex/0608013]

34. “Experimental discrimination between charge 2e/3 top quark and charge 4e/3 exotic quark production scenarios” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 98, 041801 (2007) [arXiv:hep-ex/0608044]

35. “Measurement of the ratios of the Z/G* + ¿= n jet production cross sections to the total inclusive Z/G* cross section in ppbar collisions at sqrt(s) = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] arXiv:hep-ex/0608052 FERMILAB-PUB-06-283-E(2006) (Submitted to Phys.Rev.Lett.)

36. “Search for scalar leptoquarks in the acoplanar jet topology in pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 640, 230 (2006) [arXiv:hep-ex/0607009]

37. “Search for the Standard Model Higgs Boson in the pp̄ → ZH → νν̄bb̄ channel” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 161803 (2006) [arXiv:hep-ex/0607022]

38. “Search for neutral, long-lived particles decaying into two muons in pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 161802 (2006) [arXiv:hep-ex/0607028]

39. “Search for associated Higgs boson production WH → WWW* → ℓ^{±}νℓ′+- ν′ + X in pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 151804 (2006) [arXiv:hep-ex/0607032]

40. “Search for W′ boson production in the top quark decay channel”

V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 641, 423 (2006) [arXiv:hep-ex/0607102]

41. “Search for a heavy resonance decaying into a Z+ jet final state in pp̄ collisions at \sqrt{s} = 1.96-TeV using the D0 detector” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 011104 (2006) [arXiv:hep-ex/0606018]

42. “Search for R-parity violating supersymmetry via the LL anti-E couplings λ_{121}, λ_{122} or λ_{133} in pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 638, 441 (2006) [arXiv:hep-ex/0605005]

43. “Search for neutral Higgs bosons decaying to τ pairs in pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 121802 (2006) [arXiv:hep-ex/0605009]

44. “Search for resonant second generation slepton production at the Tevatron” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 111801 (2006) [arXiv:hep-ex/0605010]

45. “Search for particles decaying into a Z boson and a photon in pp̄ collisions at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 641, 415 (2006) [arXiv:hep-ex/0605064]

46. “Search for the rare decay B_{s}^{0} → ϕμ^{+}μ^{−} with the D0 detector”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 74, 031107 (2006) [arXiv:hep-ex/0604015]

47. “Multivariate searches for single top quark production with the D0 detector” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 75, 092007 (2007) [arXiv:hep-ex/0604020]

48. “Search for squarks and gluinos in events with jets and missing transverse energy in pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 638, 119 (2006) [arXiv:hep-ex/0604029]

49. “Search for excited muons in pp̄ collisions at \sqrt{s} = 1.96- TeV”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 73, 111102 (2006) [arXiv:hep-ex/0604040] 50. “A Precise measurement of the B_{s}^{0} lifetime”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 241801 (2006) [arXiv:hep-ex/0604046]

51. “Measurement of B(t → W b)/B(t → W q) at \sqrt{s} = 1.96-TeV”

V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 639, 616 (2006) [arXiv:hep-ex/0603002]

52. “First direct two-sided bound on the B_{s}^{0} oscillation frequency”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 97, 021802 (2006) [arXiv:hep-ex/0603029]

53. “Search for pair production of second generation scalar leptoquarks in p anti-p collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 636, 183 (2006) [arXiv:hep-ex/0601047]

54. “Measurement of the isolated photon cross section in pp̄ collisions at \sqrt{s} = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 639, 151 (2006) [arXiv:hep-ex/0511054]

55. “Search for the Higgs boson in H –¿ W W(*) decays in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 96, 011801 (2006) [arXiv:hep-ex/0508054]

56. “Measurement of semileptonic branching fractions of B mesons to narrow D** states” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 171803 (2005) [arXiv:hep-ex/0507046]

57. “Measurement of the lifetime difference in the B/s0 system”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 171801 (2005) [arXiv:hep-ex/0507084]

58. “The upgraded D0 detector”

V. M. Abazov et al. [D0 Collaboration] Nucl. Instrum. Meth. A 565, 463 (2006) [arXiv:physics/0507191]

59. “Search for large extra spatial dimensions in dimuon production at D0”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 161602 (2005) [arXiv:hep-ex/0506063]

60. “Search for Randall-Sundrum gravitons in dilepton and diphoton final states” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 091801 (2005) [arXiv:hep-ex/0505018]

61. “Measurement of the W boson helicity in top quark decays”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 72, 011104 (2005) [arXiv:hep-ex/0505031]

62. “Search for single top quark production in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 622, 265 (2005) [arXiv:hep-ex/0505063]

63. “Measurement of the t anti-t production cross section in p anti-p collisions at s**(1/2) = 1.96-TeV in dilepton final states” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 626, 55 (2005) [arXiv:hep-ex/0505082]

64. “Search for neutral supersymmetric Higgs bosons in multijet events at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 151801 (2005) [arXiv:hep-ex/0504018] 65. “Production of W Z events in p anti-p collisions at s**(1/2) = 1.96-TeV and limits on

anomalous W W Z couplings” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 141802 (2005) [arXiv:hep-ex/0504019]

66. “Search for supersymmetry via associated production of charginos and neutralinos in final states with three leptons” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 151805 (2005) [arXiv:hep-ex/0504032]

67. “Measurement of the t anti-t production cross section in p anti-p collisions at s**(1/2) = 1.96-TeV using kinematic characteristics of lepton + jets events” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 626, 45 (2005) [arXiv:hep-ex/0504043]

68. “Measurement of the t anti-t production cross section in p anti-p collisions at s**(1/2) = 1.96-TeV using lepton + jets events with lifetime b-tagging” V. M. Abazov et al. [D0 Collaboration] Phys. Lett. B 626, 35 (2005) [arXiv:hep-ex/0504058]

69. “Measurement of the p anti-p –¿ W gamma + X cross section at s**(1/2) = 1.96-TeV and W W gamma anomalous coupling limits” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 71, 091108 (2005) [arXiv:hep-ex/0503048]

70. “Measurement of inclusive differential cross sections for Upsilon(1S) production in p antip collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 232001 (2005) [arXiv:hep-ex/0502030]

71. “Study of Z gamma events and limits on anomalous Z Z gamma and Z gamma gamma couplings in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 95, 051802 (2005) [arXiv:hep-ex/0502036]

72. “First measurement of sigma(p anti-p –¿ Z) x Br(Z –¿ tau tau) at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 71, 072004 (2005) [arXiv:hep-ex/0412020]

73. “Search for first-generation scalar leptoquarks in p anti-p collisions at s**(1/2) = 1.96-TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. D 71, 071104 (2005) [arXiv:hep-ex/0412029]

74. “A search for anomalous heavy-flavor quark production in association with W bosons” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 152002 (2005) [arXiv:hep-ex/0411084]

75. “A search for the flavor-changing neutral current decay B_{s}^{0} → μ^{+}μ^{−} in pp̄ collisions at \sqrt{s} = 1.96 TeV with the DØ detector” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 071802 (2005) [arXiv:hep-ex/0410039]

76. “Measurement of the ratio of B^{+} and B^{0} meson lifetimes”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 182001 (2005) [arXiv:hep-ex/0410052]

77. “Measurement of the Λ_{b}^{0} lifetime in the decay Λ_{b}^{0} → J/ψΛ^{0} with the DØ detector” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 102001 (2005) [arXiv:hep-ex/0410054] 78. “A search for W bb̄ and W H production in pp̄ collisions at \sqrt{s} = 1.96 TeV”

V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 091802 (2005) [arXiv:hep-ex/0410062]

79. “Measurement of the W W production cross section in pp̄ collisions at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 151801 (2005) [arXiv:hep-ex/0410066]

80. “A measurement of the ratio of inclusive cross sections σ(pp → Z + b − jet)/σ(pp → Z + jet) at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 161801 (2005) [arXiv:hep-ex/0410078]

81. “Measurement of dijet azimuthal decorrelations at central rapidities in pp̄ collisions at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 221801 (2005) [arXiv:hep-ex/0409040]

82. “Measurement of the B_{s}^{0} lifetime in the exclusive decay channel B_{s}^{0} → J/ψϕ” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 042001 (2005) [arXiv:hep-ex/0409043]

83. “Search for supersymmetry with gauge-mediated breaking in diphoton events at DØ” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 94, 041801 (2005) [arXiv:hep-ex/0408146]

84. “Observation and properties of the X(3872) decaying to J/ψπ^{+}π^{−} in pp̄ collisions at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 93, 162002 (2004) [arXiv:hep-ex/0405004]

85. “Search for doubly-charged Higgs boson pair production in the decay to μ^{+}μ^{+}μ^{−}μ^{−} in pp̄ collisions at \sqrt{s} = 1.96 TeV” V. M. Abazov et al. [D0 Collaboration] Phys. Rev. Lett. 93, 141801 (2004) [arXiv:hep-ex/0404015]

86. “The DZERO level 3 data acquisition system”

R. D. Angstadt et al. IEEE Trans. Nucl. Sci. 51, 445 (2004) Prepared for 13th IEEE - NPSS Real Time Conference (RT 2003), Montreal, Canada, 18-23 May 2003

87. “The DZERO level 3 data acquisition system”

D. Chapin et al. In the Proceedings of 2003 Conference for Computing in High-Energy and Nuclear Physics (CHEP 03), La Jolla, California, 24-28 Mar 2003, pp MOGT002 [arXiv:physics/0307070] FERMILAB-CONF-03-462-E(2003) Talk given at 2003 Conference for Computing in High-Energy and Nuclear Physics (CHEP 03), La Jolla, California, 24-28 Mar 2003

88. “D0 online monitoring and automatic DAQ recovery”

A. Haas et al. In the Proceedings of 2003 Conference for Computing in High-Energy and Nuclear Physics (CHEP 03), La Jolla, California, 24-28 Mar 2003, pp THGT004 [arXiv:physics/0306195] FERMILAB-CONF-03-467(2003) Talk given at 2003 Conference for Computing in High-Energy and Nuclear Physics (CHEP 03), La Jolla, California, 24-28 Mar 2003

89. “Simultaneous tracking and vertexing with elastic templates”

A. Haas Prepared for 7th International Workshop on Advanced Computing and Analysis Techniques in Physics Research (ACAT 2000), Batavia, Illinois, 16-20 Oct 2000